2.   A tank ( not rigid) contains 2.5 litres of helium gas at 280C . At constant pressure, what will be the volume of the tank after heating it and its content to 430C temperature?


1
Expert's answer
2022-05-02T05:31:07-0400

V1 = 2.5 L

T1 = 28°C + 273.15 = 301.15 K

P1 = P2 = constant

V2 = ???

T2 = 43°C + 273.15 = 316.15 K


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (2.5 L × 316.15 K) / (301.15 K) = 2.6245 L = 2.6 L

V2 = 2.6 L


Answer: The volume of the tank will be 2.6 liters